Consider a competitive firm with total costs given by TC(q) = 100 + 10q + q 2 The firm faces a market price p = 50.

(a) Write expressions for total revenue TR and marginal revenue MR as functions of output q.

(b) Write expressions for average total cost ATC, average variable cost AVC, and marginal cost MC as functions of output q.

(c) For what value of output is ATC minimized?
Answer #1

TC = 100 + 109 +9 +92 Mesa CTC) da = d (100) +10.d 190 + d (92) I da da da = 10 +29 Paso A. TR = pq =sog MR = a (TR) = so d cАта - то Avc= ve = (: = toa te 2 - (oo , 1Бt 9 > tota м се tо + 29c. ATC is minimum nenen ATC=me 100 10+9 = 10 +29 =) 100+ 10 +109 +92 = 109 +2q2 2) 292-92 = 100ATC is minimum for q=10